The probability that a 22-year-old female in the U.S. will die within one year is approximately 0.00044. If an insurance company sells a one-year, $25,000 life insurance policy to such a person for $95, what is the company's expectation? (Round your answer to the nearest dollar.)
Answer #1

Expectation value) Sum of (probability x 95x1 - 0.00044x25000 =95-11 = $84 So, the companys expectation is a profit of $84